Ronnies age plus the square of Alan's age is 2240. Alan's age plus the square of Ronnie's age 1008. How old is Ronnie and Alan?
Answer by edjones(8007)

r+a^2=2240
r^2+a=1008
.
a=1008-r^2
r+(1008-r^2)^2=2240
r^4-2016r^2+r+1016064-2240=0
r^4-2016r^2+r+1013824=0
r=31
a=47
.
Ed